### v1

# A Survey on Benchmarks of Multimodal Large Language Models

Abstract
Multimodal Large Language Models (MLLMs) are gaining increasing popularity in both academia and industry due to their remarkable performance in various applications such as visual question answering, visual perception, understanding, and reasoning. Over the past few years, significant efforts have been made to examine MLLMs from multiple perspectives. This paper presents a comprehensive review of 180 benchmarks and evaluation for MLLMs, focusing on (1)perception and understanding, (2)cognition and reasoning, (3)specific domains, (4)key capabilities, and (5)other modalities. Finally, we discuss the limitations of the current evaluation methods for MLLMs and explore promising future directions. Our key argument is that evaluation should be regarded as a crucial discipline to better support the development of MLLMs. For more details, please visit our GitHub repository: https://github.com/swordlidev/Evaluation-Multimodal-LLMs-Survey.
11footnotetext: Equal contribution: Jian Li and Weiheng Lu.
1 Introduction
Large Language Models (LLM) have recently garnered substantial interest across academic and industrial domains. The impressive performance of LLMs such as GPT ouyang2022instructGPT has fueled optimism that they could represent a step towards Artificial General Intelligence (AGI) in this era. The remarkable abilities of LLM have inspired efforts to integrate them with other modality-based models to enhance multimodal competencies. Consequently, Multimodal Large Language Models (MLLMs) jin2024emllm have emerged, This concept is further supported by the extraordinary success of proprietary models like OpenAI’s GPT-4V achiam2023gpt4 and Google’s Geminiteam2023gemini . In contrast to earlier models that were limited to solving specific tasks, Multimodal Large Language Models (MLLMs) demonstrate exceptional performance across a variety of applications, including both general Visual Question Answering (VQA) tasks and domain-specific challenges.
A comprehensive and objective benchmark for evaluating MLLMs is essential for comparing and investigating the performance of various models, and it plays a crucial role in the success of MLLMs. First, evaluating LLMs helps us better understand the strengths and weaknesses of MLLMs. For example, the SEED-Bench li2023SEED-Bench illustrates that current MLLMs show weaker abilities in understanding spatial relationships between objects while achieving relatively high performance on global image comprehension. Second, evaluations across various scenarios can offer valuable guidance for MLLM applications in fields such as medicine, industry, and autonomous driving. This, in turn, can inspire future designs and expand the scope of their capabilities. Third, the broad applicability of MLLMs underscores the importance of ensuring their robustness, safety, and reliability, especially in safety-sensitive sectors. Finally, it is significant to evaluate other user-friendly features of MLLMs including the ability to handle long contexts and accurately follow instructions. Therefore, we aim to raise awareness in the community of the importance of MLLM evaluations by reviewing the current evaluation protocols.
Recently, numerous research efforts have focused on evaluating MLLMs from various perspectives, covering factors such as perception, understanding, cognition, and reasoning. Additionally, other MLLM capabilities have been tested, including robustness, trustworthiness, specialized applications, and different modalities. Despite these efforts, a comprehensive overview that captures the full scope of these evaluations is still lacking.
Figure 1: (left) Taxonomy of this benchmarks survey, Our survey covers 5 key domains and 20-30 sub-class types, (Right up) Trend of MLLMs evaluation papers over time, (Right down) The statistics on the number of the top3 MLLMs on 83 benchmarks since 2024.
In this survey, we aim to provide a comprehensive overview of recent advancements in the rapidly evolving field of MLLM evaluation. As illustrated in Figure 1, our survey covers five key areas of MLLM evaluation, encompassing 20-30 detailed categories. The figure also illustrates the trend of MLLM evaluation papers over time, revealing a rapid increase in the number of publications. This growth indicates that the research area has garnered widespread attention. Furthermore, we provide statistics on the performance of the top three MLLMs across 83 benchmarks since 2024. The data indicates that OpenAI’s GPT-4 and Google’s Gemini exhibit superior performance and have attracted significant academic attention. as depicted in Figure. LABEL:fig:_MLLMs_evaluations_structure, we survey 180 benchmarks and organize the literature in a taxonomy consisting of five primary categories, encompassing various aspects, focusing on perception and understanding; cognition and reasoning; specific domains; key capabilities and other modalities.
  • Perception and Understanding refer to the ability to receive and extract features from multimodal data and perform cross-modal analysis. Evaluating MLLMs’ perception and understanding capabilities includes assessing whether MLLMs can perceive visual representations, identify visual details, grasp the meaning and emotions conveyed by images, and respond to related questions correctly. These abilities are the cornerstone of MLLMs, enabling them to perform a wide range of tasks and applications.
  • Cognition and Reasoning encompass the model’s capacity for advanced processing and complex inference beyond basic perception and understanding. Cognitive abilities involve processing and manipulating information to transform it into knowledge, while reasoning abilities focus on drawing logical conclusions and solving problems. Strong cognitive and reasoning abilities enable MLLMs to perform effective logical inference in complex tasks.
  • Specific domains focus on MLLMs’ capabilities in particular tasks and applications, such as their ability to process text-rich visual information and perform agent-based decision-making tasks in real-world scenarios. The discussion then extends to the evaluation of their performance in specialized domains such as medicine, autonomous driving, and industry.
  • Key capabilities significantly impact the performance and user experience of MLLMs, including managing complex dialogues, accurately following instructions, and avoiding hallucinations while maintaining trustworthiness. These capabilities are essential for ensuring that MLLMs operate effectively across diverse real-world applications and adapt to various practical scenarios.
  • Other modalities include video, audio, and 3D point clouds, which also contain rich and diverse information reflective of the real world. These modalities provide critical context and enhance the ability of MLLMs to understand complex scenarios. Evaluating MLLMs’ ability to handle various modalities can help in understanding their performance across different types of data and tasks, ensuring they are suitable for complex real-world scenarios and challenging tasks.
2 Preliminaries
Figure. 1 compares several common MLLMs including GPT4achiam2023gpt4 , Geminiteam2023gemini , LLaVAliu2023llava , Qwen-VLbai2023Qwen-VL , Claudeanthropic2024claude , InstructBLIPdai2024instructblip , mPLUG-Owl2ye2023mplugowl2 , SPHINXlin2023SPHINX , Intern-VLchen2023internvl , Yi-VLai2024Yi-VL , VideoChat2li2023videochat , Video-LLaMAzhang2023Video-LLaMA , Cambrian-1tong2024Cambrian , PLLaVAxu2024PLLaVA , Blip2li2023blip2 , MiniGPT4-Videoataallah2024MiniGPT4-Video . The standard MLLM framework can be divided into three main modules: a visual encoder $g$ tasked with receiving and processing visual inputs, a pre-trained language model that manages the received multimodal signals and performs reasoning, and a visual-language projector $P$ which functions as a bridge to align the two modalities. A diagram of the architecture and training process is illustrated in Figure. 2. This figure outlines the base LLM, the vision encoder, the projector, as well as the pertaining and instruction tuning.
2.1 MLLM Architecture
Vision Encoder
Taking the input image $X_{v}$ as input, the vision encoder compresses the original image into more compact patch features $Z_{v}$, as represented by the following formula:
$Z_{v}=g(X_{v}).$ (1)
The encoder $g$ can also be an audio encoder for audio feature extraction or an encoder for other modalities. The common vision encoders are CLIP radford2021clip , SigLIP zhai2023siglip and DINO caron2021dinov1 ; oquab2023dinov2 .
Vision-Language Projector
The task of the vision-language projector is to map the visual patch embeddings $Z_{v}$ into the text feature space:
$H_{v}=P(Z_{v}),$ (2)
where $H_{v}$ denotes the projected visual embeddings. The aligned visual features are used as prompts and inputted into the language model along with the text embeddings. Several works, such as Qformer in BLIP-2 li2023blip2 , design new projectors to reduce the number of visual tokens for efficiency.
Large Language Model
The pre-trained Large language model serves as the core component of MLLMs, endowing it with many outstanding capabilities, such as zero-shot generalization, instruction following, and in-context learning. The LLM accepts input sequences containing multiple modalities and outputs corresponding text sequences. A text tokenizer is typically bundled with the LLM, mapping text prompts $X_{q}$ to the text tokens $H_{q}$. The text tokens $H_{q}$ and the visual tokens $H_{v}$ are concatenated as the input of the language model, which outputs the final response sequence $Y_{a}$ in an autoregressive manner:
$p(Y_{a}|H_{v},H_{q})=\prod^{L}_{i=1}p(y_{i}|H_{v},H_{q},y_{<i}),$ (3)
where $L$ denotes the length of $Y_{a}$. The parameter sizes of large language models (LLMs) range from 3 billion to tens of billions. Commonly used open-source LLMs include the Llama seriestouvron2023llama ; touvron2023llama2 ; chiang2023vicuna ; zhang2024tinyllama , Phi li2023phi ; abdin2024phi3 ,Gemma gemmateam2024gemma ,Qwen bai2023qwen .
Figure 2: The architectures and training process of MLLMs.
2.2 MLLM Training
The standard training process of MLLMs is a crucial factor that determines their performance on downstream tasks and their ability to handle diverse tasks. In this section, we provide an overview of various training methodologies, including pre-training, and instruction-tuning.
Pre-training
The pre-training(PT) stage focuses on aligning different modalities within the embedding space, enabling the language model to accept inputs from various modalities. This phase primarily involves large-scale text-paired data, predominantly in the form of image-caption pairs. An image-caption pair $(X,Y)$ is typically expanded into a single-turn conversation $(X_{instruct},X_{a})$, where $X_{instruct}$ contains an image $X_{v}$ and a randomly sampled question $X_{q}$ from a set of instructions asking the assistant to briefly describe the image, and $X_{a}$ is the original image description. Given such a conversation, the model is trained to autoregressively predict the image description. Consequently, following the "next token prediction" paradigm, we can compute the probability of predicting $X_{a}$ conditioned by $X_{v}$ and optimize it using a standard cross-entropy loss function:
$\mathop{\max}_{\theta}\sum_{i=1}^{L}\log p_{\theta}(x_{i}|X_{v},X_{instruct},X_{a,<i}),$ (4)
where $L$ is the length of $X_{a}$ and $\theta$ denotes the trainable parameters. In order to better align different modalities of knowledge and avoid catastrophic forgetting during the PT stage, $\theta$ typically includes only a learnable modality interface, i.e., a vision-language projector.
Instruction-tuning
Instruction-tuning(IT) aims to fine-tune the models on specific tasks by leveraging task-specific instructions. IT is typically conducted within the paradigm of Supervised Fine-Tuning (SFT). The IT datasets are transformed into an instruction-based format, presented in the form of single-turn or multi-turn dialogue structures. Given an image $X_{v}$ and its caption, a conversation data $(X_{q}^{1},X_{a}^{1},\ldots,X_{q}^{T},X_{a}^{T})$ can be generated, where T is the total number of turns. Typically, we can organize the data into a sequence of instructions and responses, as outlined in liu2023llava . With this multimodal instruction-following sequence, IT can be performed using the same auto-regressive training objective as in the PT stage. A common strategy involves maintaining the visual encoder weights fixed while continuing to update the PT weights of both the projector and the LLM during the IT process.
3 Perception and Understanding
When evaluating the perception and understanding capabilities of MLLMs, we focus on benchmarks that assess the model’s fundamental abilities in visual information processing. This includes evaluating the MLLMs’ accuracy in object identification and detection, understanding of scene context and object relationships, and ability to respond to questions about image content. Perception and understanding abilities are the cornerstone of MLLMs, enabling them to perform a wide range of tasks and applications. This section first introduces comprehensive evaluation benchmarks for MLLMs, and then separately discusses coarse-grained and fine-grained benchmarks for visual perception.
3.1 Comprehensive Evaluation
MLLMs rely on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in various studies. In order to fully match the flourish of MLLMs, Many comprehensive evaluation benchmarks are proposed.
To advance research on visual-related tasks, LLaVA-Bench liu2023LLaVA-Bench and OwlEval ye2024OwlEval were constructed to examine a variety of MLLM capabilities, but the quantity of these benchmarks was too small to fully reflect the performance of MLLMs. There was still an issue of lacking a more comprehensive evaluation with large-scale data. Fu et al. MME filled this gap by presenting the first comprehensive MLLM evaluation benchmark, MME, which measured both perception and cognition abilities across a range of subtasks. Considering that subjective benchmarks like OwlEval relied on human labor for evaluations, which was not scalable and could exhibit significant bias, Liu et al. liu2024MMBench proposed employing GPT-4 achiam2023gpt4 to match MLLM predictions and devised MMBench to robustly assess various abilities of MLLMs. However, both the binary judgments used in MME and the multiple-choice answer selection employed by MMBench could not fully capture the complexity of open-ended real-world dialogues. In light of this concern, Open-VQA zeng2023Open-VQA and TouchStone bai2023TouchStone were proposed to support open-ended answers. Nevertheless, the relatively small scale of these benchmarks introduced instability in the evaluation statistics. To overcome this limitation, SEED-Bench was introduced with annotations six times larger than existing benchmarks. It included a substantial number of multiple-choice questions covering various evaluation dimensions for both image and video modalities. Additionally, SEED-Bench-2 categorized MLLMs’ capabilities into hierarchical levels from L0 to L4 and served as a benchmark for evaluating MLLMs’ hierarchical capabilities. Despite the promising qualitative results regarding MLLMs’ capabilities, it remained unclear how to systematically evaluate complex multimodal tasks and what the relationships among the evaluated tasks were. Based on this consideration, MM-Vet yu2023MM-Vet was devised to study integrated vision-language capabilities, allowing the evaluation to provide insights beyond overall model rankings. Moreover, MDVP-Bench lin2024MDVP-Bench and LAMM yin2023LAMM were created to provide a comprehensive assessment of MLLMs’ capabilities, particularly in understanding visual prompting instructions. Furthermore, to provide a fair and systematic assessment of MLLMs’ performance across diverse multimodal tasks, ChEF shi2023ChEF and UniBench altahan2024UniBench were constructed as standardized and holistic evaluation frameworks, which helped to comprehensively understand the capabilities and limitations of MLLMs.
3.2 Fine-grained Perception
One indispensable cornerstone of MLLMs is the ability to perceive visible objects within scenes precisely. This includes evaluating MLLMs’ capabilities in object detection and recognition, understanding details within local regions, and achieving accurate vision-language alignment. Such fine-grained perception is crucial for effective multimodal understanding and interaction.
Visual Grounding and Object Detection: Object-level grounding and detection are critical steps in better perceiving images and solving image-related questions, providing a stronger link between QA pairs and images. Flickr30k Entities plummer2016Flickr30kEntities and Visual7W zhu2016Visual7W were early benchmarks for entity localization, focusing on the detailed grounding of specific phrases in image regions and QA tasks related to local areas. These benchmarks served as the foundation for subsequent developments in the field. However, they did not provide enough detailed contextual information about objects within the scene. Zang et al. zang2023CODE investigated contextual object detection, which involved understanding visible objects within human-AI interactive contexts, and presented the CODE benchmark to facilitate research in this area. To evaluate MLLMs’ ability in challenging scenarios where images contained abundant and complex information, Wu et al. wu2023V*Bench introduced V*Bench, a benchmark focused on detailed visual grounding in high-resolution images.
Fine-grained Identification and Recognition:Fine-grained identification and recognition requires MLLMs to identify and analyze detailed visual features. In order to determine what makes for a good visual tokenizer, GVT-benchwang2023GVT-Bench was designed to evaluate MLLMs’ fine-grained visual perception through object counting and multi-class identification. However, the data used in GVT-bench were constrained to relatively small resolutions. To address this limitation, Li et al. li2023MagnifierBench extended the input resolution to a much larger range and introduced MagnifierBench, which tested MLLMs’ ability to discern the details of small objects in high-resolution images. Considering that most open-source MLLMs adopted the Pretrained Contrastive Language-Image PreTraining (CLIP) modelradford2021clip as the visual encoder, Tong et al. tong2024MMVP proposed the MMVP benchmark consisting of CLIP-blind pairs. They found that all tested models struggled with simple questions about visual details. Given that existing benchmarks were insufficient in size and did not cover crucial visual elements such as depth and spatial awareness, CV-Bench tong2024CV-Bench was created to evaluate the fundamental 2D and 3D visual understanding of MLLMs, using a significantly larger number of examples. Additionally, P2GB chen2024P2GB was designed to assess fine-grained visual capabilities, especially in text-rich scenarios, requiring comprehensive recognition and understanding of image text content. Moreover, VisualCoT shao2024VisualCoT was constructed with visual chain-of-thought prompts, allowing MLLMs to focus on specific regions within a complete image, similar to how humans comprehend intricate visual information.
Nuanced Vision-language Alignment:Nuanced vision-language alignment involves interpreting complex interactions between visual and textual information, grasping subtle meanings, and aligning semantics between images and text. Winoground thrush2022Winoground was designed to require models to match two images with two captions that contained the same set of words in different orders. However, the scale of Winoground was restricted by its costly curation, and it lacked a focus on linguistic phenomena. To address these limitations, VALSE VALSE and VLChecklist zhao2023VL-CheckList examined how MLLMs understood visual-linguistic concepts by converting real captions into confusing alternatives. By modifying textual representations related to relationships, attributes, and order, ARO yuksekgonul2023ARO assessed whether MLLMs could achieve fine-grained visual-language alignment on key concepts. Moreover, Eqbenwang2023Eqben assessed whether MLLMs were sensitive to visual semantic changes by making minimal semantic changes in images, but the image diversity was limited by virtual engines. To produce diverse images that met the requirements, Peng et al.peng2024SPEC developed a benchmark called SPEC, utilizing a progressive pipeline to synthesize images that varied in a specific attribute while ensuring consistency in other aspects.
3.3 Image Understanding
The image understanding task involves analyzing visual content to extract meaningful information, which includes grasping the context of scenes and integrating visual details with textual information to generate coherent descriptions and insights.
Multi-image Understanding: Multi-image understanding requires MLLMs to compare, analyze, and comprehend the relationships or variations among multiple images, thereby enabling more comprehensive insights into visual content. Mementoswang2024Mementos was designed to assess MLLMs’ sequential image understanding abilities, but it primarily focused on state changes among images, while neglecting other aspects. To address this issue, MileBench song2024MileBench and MuirBench wang2024MuirBench were constructed to assess multi-image cognition through a variety of task types. However, these benchmarks fell short in terms of task depth and breadth. Furthermore, MMIU meng2024MMIU provided a comprehensive evaluation by including a large number of test samples that spanned a wide range of multi-image tasks and relationships. Besides, Kil et al.kil2024CompBench focused on tasks involving relativity and comparison between multiple visual input, then introduced COMPBENCH to evaluate the comparative understanding capabilities of MLLMs. COMPBENCH required MLLMs to answer questions based on a pair of visually or semantically related images.
Implication Understanding: Understanding the meaning of images requires not only intuitive observation but also an exploration of the human emotions and cultural contexts they reflect. II-Benchliu2024II-Bench and ImplicitAVEzou2024ImplicitAVE aimed to assess MLLMs’ higher-order perceptual abilities with visual implications. To evaluate the emotional perception capabilities of MLLMs, FABA-Bench li2024FABA-Bench was designed for facial affective behavior analysis. FABA-Bench required MLLM to recognize facial expressions and movements, which are critical to understanding an individual’s emotional states and intentions.
Image Quality and Aesthetics Perception:Image quality and aesthetics perception involves assessing image quality, perceiving visual distortions, and understanding low-level attributes such as color, lighting, composition, and style. It also relates to the aesthetics and design sense of photographs. Q-Benchwu2024Q-Bench explored the potential of MLLMs in low-level perception abilities. To highlight subtle differences or similarities that might not be evident when images were viewed in isolation, Q-Bench+zhang2024Q-Bench+ extended the evaluation of low-level perception from single images to image pairs. To better align with human aesthetics, comprehensive aesthetic evaluation benchmarks AesBenchhuang2024AesBench and UNIAAzhou2024UNIAA were constructed to systematically evaluate the aesthetic abilities of MLLMs. Besides, Lin et al.lin2024DesignProbe proposed DesignProbe to comprehensively assess design capabilities of MLLMs from both the element level and the overall design level.
4 Cognition and Reasoning
MLLMs’ cognitive and reasoning abilities encompass the model’s capacity for advanced processing and complex inference beyond basic perception and understanding. Cognitive abilities involve integrating and manipulating extracted information to form coherent representations, while reasoning abilities focus on drawing logical conclusions and solving problems. Strong cognitive and reasoning abilities enable MLLMs to perform effective logical inference in complex tasks.
4.1 General Reasoning
The reasoning ability of MLLMs involves extracting and inferring relevant information from visual and textual inputs to draw logical conclusions and answer questions. This section introduces benchmarks for evaluating MLLMs’ general reasoning capabilities, focusing on three key areas: visual relation reasoning, vision-indispensable reasoning, and context-related reasoning.
Visual Relation: Evaluating MLLMs’ reasoning of visual spatial relationships involves assessing their abilities to understand the spatial arrangement, relative positions, and interactions of objects. VSRVSR and What’s Upkamath2023what'sup were developed to test MLLMs’ ability to reason about spatial relations in natural image-text pairs. However, these benchmarks primarily focused on classification tasks rather than comprehending relations within scenes. To assess more complex and general relational reasoning of MLLMs, Wang et al.wang2024CRPE introduced the Circular-based Relation Probing Evaluation (CRPE), the first benchmark to encompass all elements of relation triplets (subject, predicate, object). Recognizing that interactions and associations between distinct objects remained a significant challenge for MLLMs, Nie et al.nie2024MMRel developed Multi-Modal Relation Understanding (MMRel), a large-scale benchmark consisting of data with well-defined inter-object relations. Additionally, GSR-BENCHrajabi2024GSR-BENCH extended the What’s Up benchmark by incorporating bounding box annotations and depth information to better evaluate grounded spatial relation understanding. For the evaluation of 3D spatial cognition in MLLMs, Chen et al.cheng2024SpatialRGPT proposed SpatialRGBT-Bench, a benchmark that incorporated ground-truth 3D annotations and flexibly integrated depth information.
Context-related reasoning: Context-related visual comprehension requires MLLMs to effectively leverage contextual knowledge in solving visual problems. To assess the ability of MLLMs to answer context-dependent questions, CODIS luo2024CODIS was created, requiring MLLMs to use context from free-form text to enhance visual comprehension. Recognizing that existing MLLMs often trusted what they saw but struggled to understand presuppositions in sentences, Li et al.li2024CFMM introduced CFMM, a counterfactual reasoning benchmark designed to assess MLLMs’ ability to make presuppositions based on established facts. Zong et al. zong2024VL-ICLBench developed VL-ICL Bench, a benchmark suite designed to specifically evaluate VLLMs’ in-context learning, which involved utilizing contextual information to complete new tasks.
Vision-Indispensable reasoning: Recognizing that MLLMs may rely on language priors rather than visual information when answering questions, some works have aimed to compel MLLMs to prioritize visual data. Goyal et al.goyal2017VQAv2 introduced VQAv2, which consisted of pairs of similar images that led to different answers. However, this approach did not effectively handle open-ended questions. In response, CLEVRjohnson2016CLEVR was designed with open-ended question answering. It also ensured that external information sources, such as commonsense knowledge, did not influence answer accuracy. Nonetheless, CLEVR’s reliance on synthetic images overlooked the realism and diversity found in natural photographs. To address these limitations, GQAhudson2019GQA was developed, offering well-defined semantic representations along with the rich semantic and visual complexity of real-world images. Additionally, Chen et al.chen2024MMStar introduced MMStar, a vision-indispensable benchmark that covered a wide range of tasks and difficulty levels.
4.2 Knowledge-based Reasoning
Evaluating MLLMs’ ability to utilize knowledge is crucial for ensuring their effectiveness in complex tasks and enhancing their real-world performance. These benchmarks mainly focus on two key aspects. One aspect is knowledge-based question answering, which tests MLLMs’ ability to handle questions that require structured or extensive external knowledge. The other aspect is knowledge editing, which assesses MLLMs’ accuracy and consistency in updating and maintaining knowledge content.
Knowledge-based Visual Question Answering: Knowledge-based visual question answering requires various types of knowledge, including explicit fact-based information from knowledge bases, commonsense knowledge about human behavior, and general external knowledge. The earliest benchmarks in this area were KB-VQAwang2015KBVQA and FVQAwang2017FVQA , but they were limited by their use of "closed" knowledge sources. OK-VQAmarino2019OK-VQA advanced these benchmarks by providing a larger scale and higher quality of questions and images, utilizing "open domain" knowledge rather than a fixed source. However, OK-VQA still relied on simple lookup knowledge and required minimal reasoning. To overcome this limitation, A-OKVQAschwenk2022A-OKVQA was introduced, incorporating more commonsense knowledge and demanding greater reasoning. Furthermore, Wang et al.wang2024SOK-Bench developed SOK-Bench to evaluate MLLMs’ situated and open-world commonsense reasoning in videos.
Knowledge Editing: Knowledge editing refers to the ability to update outdated, unknown, or incorrect information within MLLMs. The benchmark MMEdit, proposed by Cheng et al.cheng2024MMEdit , provided a platform for testing the editability of MLLMs. However, it primarily focused on coarse-grained knowledge, which often failed to accurately represent fine-grained entities and scenarios in the real world. To address this limitation, Chengcheng2024MMEdit introduced MIKE, a comprehensive and challenging benchmark for fine-grained multimodal entity knowledge editing. Meanwhile, VLKEB huang2024VLKEB expanded the evaluation of knowledge editing portability, demonstrating MLLMs’ ability to effectively apply edited knowledge in related contexts. Despite that, these benchmarks overlooked the organization of multimodal knowledge and lacked a precise definition of multimodal knowledge editing. To fill this gap, MC-MKEMC-MKE was developed as a benchmark to evaluate the reliability, locality, generality, and consistency of MLLMs across different editing formats.
4.3 Intelligence&Cognition:
Inspired by the development of human intelligence, some benchmarks leverage cognitive and educational theories to assess the intelligence of MLLMs. For instance, intelligence tests featuring abstraction visual reasoning and various levels of mathematical problems are used to evaluate MLLMs’ logical reasoning capabilities. Additionally, multidisciplinary questions from various educational periods are employed to assess MLLMs’ ability to integrate diverse knowledge and apply complex reasoning skills to solve intricate problems. These approaches are crucial for understanding and enhancing the cognitive and problem-solving capabilities of MLLMs.
Intelligent Question Answering: Intelligent question answering aims to explore the intelligence of MLLMs through cognitive science perspectives. One key aspect is abstract visual reasoning (AVR)—the ability to discern relationships among patterns in images and predict subsequent patterns. RAVENzhang2019RAVEN tested abstract visual reasoning primarily with mathematical patterns over predefined geometric shapes, but its evaluation scope was limited as it addressed only a single task type. Therefore, MARVELjiang2024MARVEL and VCog-Benchcao2024VCog-Bench were introduced to evaluate MLLMs across multi-dimensional AVR tasks, but remained confined to AVR, neglecting other dimensions of cognition. To integrate cognitive science principles for a comprehensive understanding of MLLMs’ intelligence, Songsong2024M3GIA identified five key cognitive factors based on the well-recognized Cattell-Horn-Carroll (CHC) model and introduced M3GIA, the first comprehensive cognitive-driven benchmark designed to evaluate the general intelligence of MLLMs.
Mathematical Question Answering: Using mathematics problems to evaluate MLLMs is essential for assessing their logical reasoning capabilities, as these problems require complex reasoning, pattern recognition, and abstract thinking. Such tasks help determine if MLLMs can apply rules, discover patterns, and perform sophisticated reasoning. Geometry3Klu2021Geometry3K was introduced to evaluate capabilities in solving geometry problems, but it had a narrow focus on specific aspects of plane geometry. To address this issue, Lu et al.lu2024MathVista collected multiple datasets to construct an integrated benchmark, MathVista, which covered a range of mathematical tasks, such as functions and solid geometry. However, MathVista lacked a detailed classification of mathematical subdomains and emphasized visual abilities more than pure mathematical reasoning. Consequently, Math-Vwang2024Math-Vision and MathVerseMathVerse were developed, confining data to specific mathematical subjects and focusing on mathematical reasoning abilities. Additionally, Fan et al.fan2024NPHardEval4V proposed NPHardEval4V, a benchmark that used algorithmic problems and converted their textual descriptions into visual representations, aimed at evaluating the pure reasoning capabilities of MLLMs. Motivated by the idea that "if a model truly understands a problem, it should perform robustly across various tasks related to that problem," Zhou et al.zhou2024MATHCHECK introduced MATHCHECK-GEO, a benchmark focused on the universality of tasks and the robustness of question formulation within visual contexts. It was designed to test whether a model’s performance is consistent across different tasks related to the same problem.
Multidisciplinary Question Answering: Evaluating MLLMs using multidisciplinary questions from various educational stages assesses their ability to integrate and apply knowledge across different domains. This approach tests MLLMs’ reasoning and problem-solving skills in diverse contexts, providing a comprehensive measure of general intelligence and cognition. ScienceQAlu2022ScienceQA was a benchmark containing multimodal science questions with rich domain diversity. While it covered a range of disciplines, most of the questions were at the elementary to middle school level, which limited its depth. To address this issue, M3Examzhang2023M3Exam was proposed with a multilevel structure, featuring exams from three critical educational stages to comprehensively assess MLLMs’ proficiency at different levels. Additionally, SceMQAliang2024SceMQA focused on college entrance-level problems. Due to the importance of this stage, SceMQA comprised questions with answers accompanied by more detailed explanations. Intended to evaluate expert-level understanding, Yue et al.yue2024MMMU introduced MMMU, which included problems from college exams, quizzes, and textbooks across six common disciplines. However, these benchmarks were primarily available in English, restricting the evaluation to a single language. Therefore, CMMMUzhang2024CMMMU , CMMUhe2024CMMU , and MULTIzhu2024MULTI were created to evaluate multi-discipline and multi-type question understanding and reasoning in Chinese.
5 Specific Domains
This section focuses on MLLMs’ capabilities in specific tasks and applications, such as their ability to integrate complex visual and textual information, adapt to decision-making roles in dynamic environments, and effectively process diverse cultural and linguistic data. It then extends to discuss the practical applications of MLLMs, highlighting their impact on various sectors such as medicine, industry, and autonomous driving. By providing an overview of these benchmarks, this section aims to underscore the advancements in evaluating MLLMs’ performance and their potential to address real-world challenges across different domains.
5.1 Text-rich VQA
Evaluating MLLMs in text-rich visual question answering is crucial for understanding how well models interpret and integrate textual and visual information within images. This evaluation covers several aspects, including the accuracy of text recognition, contextual understanding, and the ability to synthesize information from both modalities. In addition to text comprehension, it also requires an understanding of layout and structure to effectively analyze multimodal documents, charts, and HTML.
Text-oriented Question Answering: Some works evaluated MLLMs’ effectiveness in text-related visual tasks, such as text recognition and scene text-centric visual question answering. Singh et al.singh2019TextVQA introduced TextVQA, which contained questions requiring the model to read and reason about the text in the image to provide answers. However, the short answers provided in TextVQA were insufficient for a comprehensive description of the image. To address this issue, TextCapssidorov2020TextCaps extended the length of sentences in the answers and involved many switches between OCR and vocabulary tokens. Despite these efforts, existing benchmarks could be time-consuming, and inaccurate annotations in some datasets made accuracy-based evaluation less precise. In response, Liu et al.liu2024OCRBench developed OCRBench to facilitate the accurate and convenient evaluation of MLLMs’ OCR capabilities. To quantitatively assess visual reasoning capabilities in text-rich and high-resolution scenarios, Chen et al.chen2024P2GB constructed a challenging benchmark, P2GB. This benchmark included comprehensive image understanding, fine-grained recognition, and image text content understanding. In order to cover a broader spectrum of text-rich scenarios, SEED-Bench-2-Plusli2024SEED-Bench-2-Plus was developed to evaluate MLLMs’ performance in comprehending text-rich visual data across a wide range of real-world scenarios.
Document-oriented Question Answering: Document-oriented question answering requires MLLMs not only to read text but also to interpret it within the layout and structure of the document. Mathew et al.mathew2021infographicvqa introduced InfographicVQA, which included questions requiring the combination of multiple cues. Although InfographicVQA showcased significant diversity in topics and designs, it still preferred using visual aids over long text passages. To address this limitation, Single Page DocVQA (SPDocVQA)mathew2021SPDocVQA was introduced as a more diverse benchmark, featuring documents of various types and origins created over several decades. However, it was built exclusively on single-page document excerpts and was limited to several domains represented in the Industry Documents Library. MP-DocVQAtito2023MPDocVQA extended this by including preceding and following pages of documents, but the additional pages often served as mere distractors. To overcome these issues and establish a more practical and enduring benchmark, DUDEvanlandeghem2023DUDE was created as a large-scale, multi-page, multidomain, multi-industry benchmark for evaluating MLLMs in document understanding. Additionally, to evaluate the ability of MLLMs to understand long multimodal documents, MM-NIAHwang2024MM-NIAH was designed to systematically assess long context capability of MLLMs using the Needle-In-A-Haystack (NIAH) test.
Chart-oriented Question Answering: Chart understanding plays a pivotal role when applying MLLMs to real-world tasks such as analyzing scientific papers or financial reports. ChartQAmasry2022ChartQA was a large-scale benchmark involving visual and logical reasoning over charts, but it had limited chart types. To scale up, Li et al.li2023SciGraphQA utilized Arxiv papers and constructed SciGraphQA with multi-turn question-answer conversations on scientific graphs. MMCliu2024MMC-Benchmark and ChartBenchxu2024ChartBench further expanded the types and tasks of chart data. However, the question design in these benchmarks remained relatively simplistic. To further validate MLLMs’ capabilities in more complex reasoning tasks involving chart data, Xia et al.xia2024ChartX proposed the benchmark ChartX for comprehensive chart understanding and reasoning. However, ChartX yielded artificial charts generated by GPT-4achiam2023gpt4 , leading to a narrow distribution. Therefore, CharXivwang2024CharXiv was proposed for evaluating MLLMs’ understanding of real-world scientific charts, including complex compositions with multiple subplots collected from arXiv. Moreover, in order to assess MLLMs’ understanding of charts from various perspectives, Fan et al.fan2024CHOPINLLM proposed CHOPINLLM, including three different levels of questions (literal, inferential, and reasoning QAs).
Html-oriented Question Answering: Web pages present a complex interplay of visual and textual information, along with interactive elements, requiring MLLMs to possess rigorous understanding abilities over hierarchical structures and contextual relationships. Web2Codeyun2024Web2Code was a benchmark with web pages based on instruction-response pairs. The responses included structured questions and answers about the webpage. However, web elements are often small, numerous, and scattered across the page, demanding fine-grained recognition and accurate spatial reasoning. To address these limitations, Liu et al.liu2024VisualWebBench introduced VisualWebBench, which assessed MLLMs at three levels: website-level, element-level, and action-level. Additionally, following the principle "What I cannot create, I do not understand," Plot2Codewu2024Plot2Code evaluated MLLMs’ ability to generate code that effectively rendered a provided image from HTML files, further showcasing their multimodal understanding and reasoning capabilities.
5.2 Decision-making Agents
Decision-making agents expect MLLMs to possess human-level planning and scheduling abilities, which are fundamental for making informed decisions and taking appropriate actions in complex environments. This capability holds significant potential for addressing real-world problems.
Embodied Decision-making: Embodied Decision-making requires MLLMs to be able to integrate sensory inputs and interact with the environment in a way that mimics human physical experiences. OpenEQAOpenEQA2023 was the first benchmark for embodied question answering, supporting both episodic memory and active exploration use cases. However, it focused solely on the answers provided by MLLMs and did not consider the intermediate reasoning processes. Chen et al.chen2023PCA-EVAL argued that it was essential to enable multi-dimensional evaluation of the decision-making process, encompassing perception, reasoning, and action perspectives, rather than relying solely on final rewards or success rates. They proposed PCA-EVAL for evaluating the embodied decision-making ability of MLLMs from different perspectives, including three complex scenarios: autonomous driving, domestic robotics, and open-world games. However, these benchmarks were limited by the small number of handcrafted questions and reliance on single-image visual observation. To study the embodied planning and decision-making capabilities of MLLMs more systematically, Chen et al.chen2024EgoPlan-Bench used large-scale egocentric videos reflecting daily human activities from a first-person perspective to construct EgoPlan-Bench. EgoPlan-Bench aimed to assess MLLMs’ human-level planning capabilities in real-world scenarios, featuring realistic tasks, diverse actions, and complex visual observations. Despite that, these benchmarks failed to sufficiently challenge or showcase the full potential of MLLMs in complex environments. To address this gap, Liu et al.liu2024VisualAgentBench introduced VisualAgentBench (VAB), a comprehensive and pioneering benchmark specifically designed to train and evaluate MLLMs as visual foundation agents across diverse scenarios, with tasks formulated to probe the depth of MLLMs’ understanding and interaction capabilities.
Mobile Agency: Mobile-Agent leverages visual perception tools to accurately identify and locate both visual and textual elements within the app’s front-end interface. Drawing on the visual information, it autonomously plans and decomposes complex tasks, navigating the mobile app through each step of the operation. To comprehensively assess Mobile-Agent’s capabilities, Wang et al.wang2024Mobile-Eval introduced Mobile-Eval, a benchmark centered around current mainstream mobile apps. Mobile-Eval included instructions for various difficulty levels, but it primarily assessed whether MLLMs could complete the instructions, neglecting the fine-grained perception of the web interface. To fill this gap, You et al.you2024Ferret-UI proposed Ferret-UI, which required MLLMs to explain the functionality and provide fine-grained descriptions of UI elements. However, existing benchmarks for MLM agents in interactive environments were constrained by single environments and lacked detailed and generalized evaluation methods. To overcome these limitations, Xu et al.xu2024CRAB introduced CRAB, the first agent benchmark framework designed to support cross-environment tasks.
5.3 Diverse Cultures and Languages
Most benchmarks primarily use English, leading to the neglect of other languages and cultures. To address this limitation, some benchmarks have been introduced to supplement data in a broader range of languages. CMMUhe2024CMMU , CMMMUzhang2024CMMMU , and MULTIzhu2024MULTI were presented for multi-modal and multi-type questions in Chinese, featuring a wider variety of question types. Besides, the Henna benchmarkalwajih2024Henna was proposed to test MLLMs in Arabic culture, while the LaVy-Bench benchmarktran2024LaVy was designed for evaluating MLLMs’ understanding of Vietnamese visual language tasks. To further enrich language diversity, MTVQAtang2024MTVQA was proposed as the first benchmark featuring high-quality human expert annotations across 9 diverse languages. However, although MTVQA extended its linguistic range, it kept images the same, resulting in a narrow cultural representation. To address this limitation, Romeroromero2024CVQA constructed CVQA, a culturally diverse multilingual visual question answering benchmark designed to cover a rich set of languages and cultures.
5.4 Other Applications
Some works focused on assessing MLLMs’ abilities to handle highly professional and domain-specific data, such as medicine, transportation, engineering, remote sensing, and autonomous driving. These evaluations provided insights into how well MLLMs could adapt to and process specialized information in various complex fields, highlighting their potential for applications in areas requiring deep expertise and precise knowledge.
Geography and Remote Sensing:MLLMs can help enhance geographic information extraction and environmental monitoring by analyzing multimodal data in the fields of geography and remote sensing. Roberts et al.roberts2024ChartingNewTerritories constructed a benchmark to probe the key geographic and geospatial knowledge of a suite MLLMs. However, the diverse geographical landscapes and varied objects in remote sensing imagery were not adequately considered. To bridge this gap, Muhtar et al.muhtar2024LHRS-Bench introduced LHRS-Bench, a benchmark for thoroughly evaluating MLLMs’ abilities in remote sensing image understanding.
Medicine:Medical-based benchmarks primarily assess MLLMs’ ability to integrate medical knowledge with visual modalities to perform accurate medical diagnoses and recommendations. The Asclepius benchmarkwang2024Asclepius covered a range of medical specialties, aiming to comprehensively evaluate MLLMs’ capabilities across various medical fields. However, it primarily focused on 2D medical images, leaving 3D images less explored. To address this, Bai et al.bai2024M3D introduced M3D-Bench, a 3D multimodal medical benchmark that enabled automatic evaluation across eight tasks. Furthermore, Chen et al.chen2024GMAI-MMBench developed GMAI-MMBench, the most comprehensive general medical AI benchmark to date, featuring a well-categorized data structure and multi-perceptual granularity.
Industry: Some benchmarks have been designed to assess MLLMs’ capabilities in industrial design and manufacturing applications. Doris et al. doris2024DesignQA proposed a benchmark called DesignQA to explore MLLMs’ understanding of design based on engineering requirement documents. DesignQA integrated information from both visual and long-text inputs, emphasizing the complexity and multimodal nature of real-world engineering tasks. To evaluate MLLMs’ abilities in robotic applications, the MMRo li2024MMRo benchmark defined four fundamental capabilities required for a robot’s central processing unit and was the first diagnostic benchmark specifically designed to systematically analyze the diverse failure modes of MLLMs in robotics.
Society: Benchmarking MLLMs in addressing social needs and related domains is essential for evaluating their performance in real-world interactions and practical scenarios. To address the needs of blind individuals, Gurari et al. gurari2018VizWiz proposed VizWiz, a benchmark consisting of visual questions originating from blind people. To evaluate MLLMs’ abilities in understanding information from social media interactions, Jin et al. jin2024MM-Soc introduced MM-SOC, a benchmark designed to holistically assess MLLMs’ performance on multimodal tasks derived from online social networks. Considering the critical role of transportation in modern society, Zhang et al. zhang2024TransportationGames developed TransportationGames, a comprehensive benchmark aimed at accurately evaluating MLLMs’ capabilities in performing transportation-related tasks.
Autonomous Driving: Autonomous driving is a rapidly developing field with immense potential to improve transportation safety and efficiency through advancements in sensor technologies and computer vision. Qian et al.qian2024NuScenes-QA constructed the first VQA benchmark for autonomous driving scenarios, named NuScenes-QA. It addressed visual question answering and provided richer visual information, including images and point clouds. However, NuScenes-QA primarily focused on scene-level driving tasks and could not cover all the reasoning processes involved in driving. In contrast, DriveLM-Datasima2024DriveLM-Data was a more comprehensive benchmark, which encapsulated perception, prediction, and planning for autonomous driving.
6 Key Capabilities
These benchmarks evaluated dialogue capabilities, including handling extended dialogues and accurately following instructions, as well as assessing the model’s level of hallucination and trustworthiness. Such capabilities are crucial for ensuring that MLLMs perform effectively across a range of real-world applications and can adapt to various practical scenarios.
6.1 Conversation Abilities:
Some benchmarks focus on evaluating MLLMs’ performance in conversations, specifically assessing how well these models handle long contexts and follow complex instructions accurately. Such evaluations are crucial for ensuring that MLLMs can engage effectively in diverse dialogues and deliver reliable performance in real-world applications.
Long-context Capabilities: Due to the window length limitations inherent in MLLMs’ architectures, evaluating their ability to handle long contexts is challenging. This involves assessing whether MLLMs can maintain accurate recall and effective understanding as the amount of contextual information increases. MileBench song2024MileBench and MMNeedle wang2024MMNeedle explored the long-context recall abilities of MLLMs using needle-in-a-haystack (NIAH) method and image retrieval tasks. However, evaluating the long-context understanding of MLLMs in videos remained a significant challenge. Therefore, Zhou et al. zhou2024MLVU proposed MLVU, which was developed using long videos of diversified lengths.
Instruction Adherence: Instruction adherence requires that MLLMs execute complicated instructions. This involved not only recognizing the content of the instructions but also meticulously executing the detailed demands without deviation. Wang et al.wang2023DEMON built a benchmark called Demon for demonstrative instruction understanding. However, it focused only on demonstrative instruction following and ignored other flexible instruction scenarios. To explore how MLLMs performed on broader, open-ended prompts, Bitton et al.bitton2023VisIT-Bench created VisIT-Bench to cover a wide array of ’instruction families’ that resembled real-world user behavior. Although these benchmarks evaluated the basic instruction-following capabilities of MLLMs, the ability of MLLMs to adapt to new instructions while incorporating both old and new ones remained unclear. Therefore, Chen et al.chen2024CoIN presented a benchmark named Continual Instruction Tuning (CoIN) to assess MLLMs in a sequential instruction tuning paradigm. To better measure MLLM adherence to instructions, MIA-Benchqian2024MIA-Bench was introduced to test how well MLLMs follow layered instructions and generate accurate responses matching specific patterns.
6.2 Hallucination
Hallucination refers to information in LVLMs’ responses that does not accurately reflect the visual input, which poses potential risks of substantial consequences.
To measure object hallucination, Rohrbach et al.rohrbach2019CHAIR proposed CHAIR (Caption Hallucination Assessment with Image Relevance), which assessed captioned objects that were actually present in an image. However, CHAIR was unstable and required complex human-crafted parsing rules for exact matching. Alternatively, POPEli2023POPE converted hallucination into a binary classification problem, but it required the input questions to follow specific templates, such as ’Is there a/an <object> in the image?’. In comparison, GAVIEliu2024GAVIE can evaluate model hallucination in an open-ended manner without requiring ground-truth answers or pre-designed instruction formats. However, it still focused on evaluating object hallucinations, neglecting other types of hallucinations in MLLMs. M-HalDetectgunjal2024M-HalDetect and MMHAL-BENCH sun2023MMHAL-BENCH extended the scope of previous works by not only considering hallucinations related to objects but also addressing other categories such as object attributes and spatial relations. Moreover, Chen et al.chen2024MHaluBench presented MHaluBench, a meta-evaluation benchmark that encompassed various hallucination categories and multimodal tasks. Additionally, Zhang et al.zhang2024MRHal-Bench introduced MRHalBench to evaluate hallucinations in multi-round dialogues. To assess MLLMs’ hallucination in video understanding, Wang et al.wang2024VideoHallucer introduced VideoHallucer, the first comprehensive benchmark for hallucination detection in videos.
Some benchmarks aimed to explore more cost-effective and feasible methods for evaluating hallucinations. Wang et al.wang2023HaELM argued that hallucinations measured using object-based evaluations like POPE merely exploited the judgment bias present in MLLMs, rather than reflecting their actual hallucinations. They proposed Hallucination Evaluation based on Large Language Models (HaELM), the first to utilize LLMs for hallucination evaluation within MLLMs. Considering that reliance on LLMs resulted in significant costs, Wang et al.wang2024AMBER proposed an LLM-free multi-dimensional benchmark, AMBER. AMBER provided comprehensive coverage of evaluations for various types of hallucinations and offered detailed annotations to support an LLM-free evaluation pipeline. Despite that, many of the visual hallucination (VH) instances in these benchmarks came from existing datasets, which resulted in a biased understanding of MLLMs’ performance due to the limited diversity of such VH instances. Therefore, Huang et al.huang2024VHTest proposed generating diverse VH instances using a text-to-image generative model. Based on this method, they collected a benchmark dataset called VHTest. Additionally, to eliminate the need for costly data annotation and minimize the risk of training data contamination, Cao et al.cao2024MMECeption proposed an annotation-free evaluation method that required only unimodal data to measure inter-modality semantic coherence and inversely assessed MLLMs’ tendency to hallucinate.
There are also some works that explored the causes and mechanisms of hallucination. Observing that MLLMs’ strong language bias often overshadowed visual information, leading to an overreliance on language priors rather than visual context, HallusionBenchguan2024HallusionBench was proposed to focus on diagnosing both the visual illusion and knowledge hallucination of MLLMs. Cui et al.cui2023Bingo constructed a benchmark called Bingo, which systematically categorized and analyzed the reasons behind the occurrence of hallucinations. Moreover, Han et al.han2024CorrelationQA identified a typical class of inputs that baffled MLLMs: images that were highly relevant but inconsistent with answers, causing MLLMs to suffer from hallucination. To quantify this effect, they proposed CorrelationQA, the first benchmark that assessed the hallucination level given spurious images.
6.3 Trustworthiness
Evaluating the trustworthiness of multimodal large language models (MLLMs) encompasses various aspects, including accuracy, consistency across different scenarios, and safety in handling sensitive content. This section focuses on benchmarks that assess MLLMs specifically in terms of robustness and safety. Robustness examines how well the model performs with diverse or unexpected inputs, ensuring reliable outputs across various conditions. Safety evaluates the model’s capacity to avoid generating harmful or inappropriate content, thereby protecting users from potential risks. These dimensions are essential for confirming that MLLMs are dependable and safe for real-world applications.
Robustness: Robustness evaluation in MLLMs involves assessing how well the models handle input disturbances and maintain consistent performance across different data distributions and tasks. Some studies tested the robustness of MLLMs by examining their performance across various input distributions. For instance, to examine MLLMs’ visual capability in terms of visual diversity, Cai et al.cai2023benchlmm proposed BenchLMM, a benchmark that assessed MLLMs in three distinct styles of distribution shifts: artistic style, sensor style, and application style. In terms of prompt diversity, MLLMs might generate erroneous responses when faced with deceptive or inappropriate instructions. To quantitatively assess this vulnerability, MMR liu2024MMR and MAD-Bench qian2024MAD-Bench were constructed to comprehensively evaluate MLLMs’ capability to resist deceiving or misleading information in the prompt. Additionally, some incorrect responses from MLLMs stem from their difficulty in understanding what they can and cannot perceive in images, a capability known as self-awareness in perception, which undermines their robustness. Therefore, MM-SAP wang2024MM-SAP and VQAv2-IDK cha2024VQAv2-IDK were developed to assess MLLMs’ understanding of self-awareness, particularly in scenarios where the correct response is ’I Don’t Know’. Ye et al.ye2024MM-SpuBench explored the issue of spurious bias, finding that MLLMs often relied on spurious correlations between non-essential input attributes and target variables for their predictions. To better understand this problem, they introduced MM-SPUBENCH, a comprehensive benchmark designed to evaluate MLLMs’ dependence on nine distinct categories of spurious correlations.
Safety: Some works primarily test the safety of MLLMs, ensuring they refrain from producing content that poses ethical risks or causes social harm. Driven by the observation that MLLMs tended to respond to malicious questions when a query-relevant image was presented in the dialogue, Liu et al.liu2024MM-SafetyBench constructed a safety-measurement benchmark called MM-SafetyBench, which encompassed a wide range of unsafe scenarios. Meanwhile, by designing specific inputs, jailbreak attacks could induce the model to generate harmful content that might violate human values. To explore this issue, JailBreakVluo2024JailBreakV-28K was developed to assess various jailbreak techniques on MLLMs. Besides, MLLMs were vulnerable to data extraction and privacy leaks during inference, posing risks in privacy-sensitive applications. Therefore, it was crucial to empower privacy information with the "right to be forgotten" through machine unlearning, which removed private or sensitive information from models. To address this issue, Li et al.li2024MMUBench established MMUBench to evaluate the efficacy, generality, specificity, fluency, and diversity of machine unlearning methods in MLLMs. Moreover, Shi et al.shi2024SHIELD introduced SHIELD, a benchmark designed to evaluate the effectiveness of MLLMs in addressing various challenges within the domain of facial security, including face anti-spoofing and face forgery detection. However, these benchmarks typically examined only one or a few aspects of safety and trustworthiness, lacking a comprehensive evaluation. To address this issue, MultiTrust zhang2024MultiTrust and RTVLM li2024RTVLM were proposed as comprehensive and unified benchmarks. Both evaluated the safety capabilities of MLLMs across diverse dimensions and tasks.
7 Other Modalities
Beyond the image modality, other modalities such as video, audio, and 3D point clouds also contain rich and diverse information reflective of the real world. These modalities provide critical context and enhance the ability to understand complex scenarios. Evaluating MLLMs across these varied modalities is crucial for developing robust and versatile models capable of handling a wide range of complex real-world tasks. This section provides an overview of benchmarks designed to assess MLLMs across these different modalities, aiming to highlight their strengths and identify areas for further improvement.
7.1 Videos
Compared to images, the video modality features temporal dynamics and sequential context, involving changes over time and interactions between frames. Therefore, evaluating MLLMs on video-related aspects, such as temporal coherence and action understanding, is crucial for assessing their performance in understanding and interpreting video content.
Temporal Perception: Temporal perception is a fundamental distinction between video-centered and image-centered applications. TimeIT Ren2023TimeIT was conducted to assess the temporal understanding capabilities of MLLMs. However, TimeIT neglected the distinction between various temporal aspects. To address this issue, MVBench li2024MVBench , Perception Test pătrăucean2023PerceptionTest , VilMA kesen2023ViLMA , and VITATECS li2023VITATECS introduced a range of fine-grained temporal aspects, enabling a more comprehensive and nuanced evaluation of temporal perception capabilities. However, MVBench, VITATECS, and VilMA were limited to single-task formats, and Perception Test was constrained to indoor videos, making these benchmarks less suitable for evaluating MLLMs. In response to these issues, Liu et al. liu2024TempCompass proposed TempCompass, a benchmark designed to evaluate the temporal perception abilities of MLLMs across various temporal aspects, different task formats, and diverse types of videos. Moreover, focusing on the causal relationships between concrete actions and their effects, Nguyen et al. nguyen2024OSCaR proposed OsCaR, a benchmark that included various tasks for understanding causal temporal states. However, the videos in these benchmarks typically had clear temporal structures and scene semantics, whereas activities of daily living often involve temporal unstructuredness, with diverse actions occurring concurrently. To address this, ADLMCQ chakraborty2024ADLMCQ was introduced to evaluate MLLMs on tasks related to daily living activities.
Long Video Understanding: Video MLLMs often use key frame extraction for video understanding tasks, which presents challenges as video length increases. Videos in most benchmarks had clear length limitations, which were insufficient to reflect MLLMs’ long-video understanding capabilities. Some works, like Egoschemamangalam2023EgoSchema and MovieChat-1ksong2024MovieChat , collected long videos and created questions based on them. MovieChat-1k was a benchmark for long video understanding tasks, but many of the questions in MovieChat-1k targeted exact time segments, which degraded the tasks to short-video problems. EgoSchema presented video reasoning tasks using first-person footage. However, it only focused on a few aspects of long videos, rather than offering a comprehensive analysis of long-video understanding. To address these issues, MLVUzhou2024MLVU was developed as a comprehensive benchmark with various task categories to evaluate MLLMs’ capabilities in understanding long videos. Additionally, Du et al.du2024Event-Bench introduced Event-Bench, an event-oriented long-video understanding benchmark that used diverse videos to evaluate MLLMs’ ability to understand complex event narratives.
Comprehensive Evaluation: Comprehensive evaluation of video MLLMs focuses on their overall abilities, including video comprehension, scene recognition, temporal reasoning, and various other related tasks. With the development of video MLLMs, several works collected extensive video datasets to comprehensively evaluate MLLMs, such as Video-Bench ning2023Video-Bench and AutoEval-Video chen2024AutoEval-Video . However, they were limited by a lack of diversity in video types and insufficient coverage of temporal dynamics. To create a more comprehensive and high-quality assessment of MLLM in video, Fu et al. fu2024Video-MME proposed Video-MME, which offered more diverse video types, broader temporal coverage, and higher quality annotations. Additionally, due to the ability of videos to capture rich representations of real-world dynamics and causalities, some works evaluated MLLMs’ understanding of videos through the lens of "world models," which required MLLMs to interpret and reason about complex real-world dynamics. MMWorld he2024MMWorld and WorldNet ge2024WorldNet-Crafted were designed to rigorously evaluate the capabilities of MLLMs in world modeling via video understanding. MMWorld spanned a broad spectrum of disciplines, featuring a diverse array of question types for multi-faceted reasoning, while WorldNet encompassed millions of samples across a wide range of real-world scenarios and tasks.
7.2 Audio
MLLMs that jointly process audio and language hold great promise for audio understanding. Dynamic-SUPERB huang2024Dynamic-SUPERB was a benchmark that assessed MLLMs’ ability to follow instructions in the audio domain. Nevertheless, Dynamic-SUPERB focused solely on human speech processing, ignoring other types of audio. To measure the understanding abilities of MLLMs in music, Weck et al. weck2024MuChoMusic presented MuChoMusic, the first benchmark for evaluating music understanding in audio MLLMs. In order to offer comprehensive coverage of audio signals, including human speech, natural sounds, and music, Yang et al. yang2024AIR-Bench presented AIR-Bench, a comprehensive benchmark designed to evaluate MLLMs’ ability to comprehend various audio signals and interact according to instructions.
7.3 3D Scenes
3D scenes offer a significant advantage over 2D images by providing accurate spatial relationships, depths, and occlusions, which are essential for interpreting complex environments. For MLLMs, enhanced 3D scene perception enables more precise responses in applications such as navigation, augmented reality, and complex spatial reasoning tasks. ScanQA azuma2022ScanQA was proposed for 3D visual question answering, requiring models to answer given questions based on point clouds. It was formatted as an attribute classification task, which made it difficult to assess MLLMs’ detailed understanding of 3D scenes. LAMMyin2023LAMM transformed the classification task into a multiple-choice problem and added two tasks beyond 3D visual question answering: 3D object detection and visual grounding.
Table 1: Summary of existing MLLM evaluations and benchmarks. This table provides an overview of the key attributes of the benchmarks introduced in this paper. The "Focus" describes the primary features and focal points of each benchmark; The "Answer Annotation" denotes the size of the dataset for each benchmark;The "Answer Type" categorizes the format of responses: T/F refers to True or False, Y/N denotes Yes or No, MQA indicates multiple-choice answers, restrictive text refers to text with specific format constraints, and open indicates open-ended responses. "Evaluation" specifies the evaluation methods employed;The "Models" column indicates the number of models evaluated in each benchmark.
Benchmark Focus Answer Annotation Answer Type HGEvaluation Models
VSRVSR visual spatial reasoning 10972 T/F N/A 4
VALSEVALSE visio-linguistic grounding capabilities 6795 A/Bs N/A 6
VL-CheckListzhao2023VL-CheckList image-text matching A/B N/A 7
AROyuksekgonul2023ARO composition and order understanding MQA N/A 4
Eqbenwang2023Eqben video-text matching 250k A/B N/A 9
LLaVA-Benchliu2023LLaVA-Bench visual instruction following 54 open GPT 4
OwlEvalye2024OwlEval comprehensive evaluation 82 open Human 5
OCRBenchliu2024OCRBench OCR capabilities 1000 open N/A 14
DUDEvanlandeghem2023DUDE understanding visually-rich documents 11448 open N/A 6
POPEli2023POPE object hallucination 3000 Y/N N/A 5
PerceptionTestpătrăucean2023PerceptionTest Memory, Abstraction, Physics, Semantics 11.6k MQA, open N/A
NuScenes-QAqian2024NuScenes-QA autonomous driving 460k open N/A 9
CODEzang2023CODE contextual object detection 10346 open - 2
M3Examzhang2023M3Exam official human exam questions 12317 MQA N/A 7
LAMMyin2023LAMM instruction tuning 62,439 open GPT -
LAMMyin2023LAMM multi-modal instruction tuning 62439 open GPT 4
MMEMME perception and cognition 1.5k Y/N N/A 10
MMBenchliu2024MMBench objective evaluation 3217 open GPT 14
MovieChatsong2024MovieChat long video understanding 14k open GPT 6
MM-Vetyu2023MM-Vet integrated capabilities 205 open GPT 9
SciGraphQAli2023SciGraphQA scientific question-answering 657k open GPT 5
DEMONwang2023DEMON demonstrative instruction understanding 477.72K open N/A 9
M-HalDetectgunjal2024M-HalDetect hallucination detection 16K open human 3
VisIT-Benchbitton2023VisIT-Bench instruction following 592 open Human/GPT 14
EgoSchemamangalam2023EgoSchema long video understanding 5063 MQA N/A 4
TouchStonebai2023TouchStone open-ended real-world dialogues 908 open GPT 7
Q-Benchwu2024Q-Bench low-level visual perception 2990 Y/N,open N/A 15
MMHAL-BENCHsun2023MMHAL-BENCH penalizing hallucinations 96 open GPT 6
PCA-EVALchen2023PCA-EVAL decision-making ability 300 MQA GPT 6
MathVistalu2024MathVista mathematical tasks 735 MQA,open GPT 12
MMEditcheng2024MMEdit knowledge editing open N/A 2
HallusionBenchguan2024HallusionBench image-context reasoning 1129 open GPT 15
What’sUpkamath2023what'sup spatial relations 820 text options N/A 8
ChEFshi2023ChEF Comprehensive Evaluation Framework - - - 9
Bingocui2023Bingo hallucinations related to bias and interference 370 open GPT 2
MagnifierBenchli2023MagnifierBench visual perception of small objects 283 MQA, open Human/GPT -
ViLMAkesen2023ViLMA visio-linguistic capabilities 4177 caption matching N/A 12
AMBERwang2024AMBER hallucinations related to attributes and relations 15k Y/N, open N/A 9
MMC-Benchmarkliu2024MMC-Benchmark visual charts understanding 2k Open,MQA GPT 6
AutoEval-Videochen2024AutoEval-Video open ended video question answering 327 open GPT 11
Video-Benchning2023Video-Bench comprehensive evaluation for video 17036 MQA N/A 8
MMMUyue2024MMMU college-level subject knowledge 11.5k MQA,open N/A 23
MVBenchli2024MVBench video understanding 4000 MQA N/A 17
VITATECSli2023VITATECS temporal concept underStanding 13k open N/A 9
MM-SafetyBenchliu2024MM-SafetyBench safety-critical evaluations 5040 open GPT 12
TimeITRen2023TimeIT long video understanding 125k open 8
EgoPlan-Benchchen2024EgoPlan-Bench human-level planning 5k MQA N/A 28
M3DBenchli2023M3DBench 3D instruction following 327k open GPT 3
V*Benchwu2023V*Bench grounding detailed visual information 191 MQA 7
DriveLM-DATAsima2024DriveLM-Data autonomous driving
ChartBenchxu2024ChartBench chart comprehension 600k open N/A 21
TransportationGameszhang2024TransportationGames transportation 1910 MQA,T/F,open GPT 16
MMVPtong2024MMVP visual perception 300 MQA 9
BenchLMMcai2023benchlmm robustness open GPT 10
MM-SAPwang2024MM-SAP self-awareness in perception for MLLMs 1150 MQA N/A 14
AesBenchhuang2024AesBench image aesthetics perception 8400 MQA GPT 15
SPECpeng2024SPEC fine-grained comprehension text options N/A 5
Mementoswang2024Mementos sequential image reasoning abilities 4761 free-from GPT 9
CMMMUzhang2024CMMMU college-level subject knowledge 12K open N/A 11
RTVLMli2024RTVLM red teaming 5200 open GPT 10
CMMUhe2024CMMU knowledge comprehension and reasoning 3603 open N/A 10
Mobile-Evalwang2024Mobile-Eval mobile device operations 33 N/A N/A
LHRS-Benchmuhtar2024LHRS-Bench remote sensing 690 MQA N/A 9
MHaluBenchchen2024MHaluBench multimodal hallucination detection 420 open N/A 2
CorrelationQAhan2024CorrelationQA instinctive bias across different types 7308 open N/A 9
SHIELDshi2024SHIELD face spoofing and forgery detection T/F, MQA N/A 2
SceMQAliang2024SceMQA core science subjects 1045 open GPT 6
Q-Bench+zhang2024Q-Bench+ low-level visual perception 949 Y/N,open N/A 24
VQAv2-IDKcha2024VQAv2-IDK self-awareness hallucination 20k open N/A 3
Asclepiuswang2024Asclepius medicine 3232 MQA N/A 6
ChartXxia2024ChartX chart reasoning 48k open N/A 11
MCUBchen2024MCUB understand inputs from diverse modalities MQA N/A 5
MAD-Benchqian2024MAD-Bench resist deceiving information in the prompt 1000 open GPT 19
CODISluo2024CODIS context-dependent visual comprehension 706 Close&Open-ended Human/ GPT 14
MM-Socjin2024MM-Soc social media content MQA,open N/A 10
VHTesthuang2024VHTest visual hallucination 1200 Y/N N/A 8
Math-Visionwang2024Math-Vision mathematical reasoning 3k open GPT 9
MIKEli2024MIKE knowledge editing 1000 entities open N/A 2
MMEcaptioncao2024MMECeption inter-modality semantic coherence N/A N/A 7
OSCaRnguyen2024OSCaR object state understanding 14084 open N/A 8
CRPEwang2024CRPE relation comprehension - MQA N/A 3
TempCompassliu2024TempCompass temporal perception 7540 multiple types of tasks GPT 11
Hennaalwajih2024Henna Arabic culture 1132 open GPT 2
NPHardEval4Vfan2024NPHardEval4V reasoning abilities 900 restrictive text N/A 9
VLKEBhuang2024VLKEB knowledge editing 3174 open N/A 5
CoINchen2024CoIN continual instruction tuning open N/A 2
VL-ICL Benchzong2024VL-ICLBench in-context learning 1520 open N/A 12
MathVerseMathVerse visual math tasks 15k open GPT 17
VisualCoTshao2024VisualCoT specific local region identification free-from GPT
P2GBchen2024P2GB visual reasoning capabilities 2130 MQA - 3
MDVP-Benchlin2024MDVP-Bench visual prompting research - Y/N GPT 4
MMStarchen2024MMStar vision-indispensable tasks 1500 MQA N/A 16
M3Dbai2024M3D 3D medical tasks MQA,open N/A 2
JailBreakV-28Kluo2024JailBreakV-28K the robustness of MLLMs against jailbreak attacks 28000 open LLM 10
FABA-Benchli2024FABA-Bench facial affective behavior analysis 403 open N/A 5
Ferret-UIyou2024Ferret-UI understanding of mobile UI screens 69 N/A N/A 4
VisualWebBenchliu2024VisualWebBench understanding and grounding in web scenarios 1.5k open N/A 14
DesignQAdoris2024DesignQA real-world engineering tasks 1451 open N/A 5
Lavytran2024LaVy vietnamese visual language understanding
UNIAAzhou2024UNIAA image aesthetic attributes 5354 multiple types of tasks N/A 13
CFMMli2024CFMM counterfactual reasoning 2400 A/B N/A 7
MARVELjiang2024MARVEL abstract visual reasoning 770 MQA N/A 9
DesignProbelin2024DesignProbe design task open GPT 9
ImplicitAVEzou2024ImplicitAVE implicit value extraction 1610 restrictive text N/A 6
MMT-Benchying2024MMT-Bench general purpose multimodal intelligence 31k MQA GPT 30
SEED-Bench-2-Plusli2024SEED-Bench-2-Plus text-rich visual comprehension 2.3k MQA N/A 34
MileBenchsong2024MileBench long-context capabilities 6440 MQA N/A 22
MileBenchsong2024MileBench multimodal long-context capabilities 6440 MQA N/A 22
Plot2Codewu2024Plot2Code multi-modal code tasks 132 open GPT 14
SOK-Benchwang2024SOK-Bench video reasoning 44K MQA N/A 7
MRHal-Benchzhang2024MRHal-Bench hallucinations in multi-round dialogues 105 open GPT 15
MTVQAtang2024MTVQA multilingual text-rich scenarios 28607 open N/A 18
MMUBenchli2024MMUBench evaluation of machine unlearning 1000 open N/A 4
Video-MMEfu2024Video-MME comprehensive evaluation in video analysis 2700 MQA N/A 12
SpatialRGPTcheng2024SpatialRGPT grounded spatial reasoning 1406 open GPT 9
MLVUzhou2024MLVU multi-task long Video Understanding 2593 MQA,open GPT 20
MLVUzhou2024MLVU multi-task long video understanding 2593 MQA N/A 20
M3GIAsong2024M3GIA intelligence tests 1800 MQA N/A 24
II-Benchliu2024II-Bench higher-order perceptual 1434 MQA N/A 20
CVQAromero2024CVQA culturally-diverse multilingual VQA 9k MQA N/A 8
MultiTrustzhang2024MultiTrust the trustworthiness of MLLMs open N/A 21
MM-NIAHwang2024MM-NIAH long multimodal documents 12k restrictive text 9
MMWorldhe2024MMWorld multi-discipline, multi-faceted video understanding 6627 MQA, open GPT 12
MMRelnie2024MMRel inter-object relations 15k Y/N N/A 6
ADLMCQchakraborty2024ADLMCQ activities of daily living MQA N/A 5
MuirBenchwang2024MuirBench multi-image understanding 2600 MQA N/A 20
VCog-Benchcao2024VCog-Bench abstract visual reasoning 1440 MQA N/A 16
MMRliu2024MMR robustness to leading questions 600 MQA N/A 18
MMNeedlewang2024MMNeedle long-context capabilities 280K restrictive text N/A 11
MC-MKEMC-MKE knowledge editing 2884 open N/A 2
GSR-BENCHrajabi2024GSR-BENCH spatial relationships 820 Template-based generation N/A 9
Event-Benchdu2024Event-Bench event-oriented long video understanding 2190 MQA N/A 12
VideoHallucerwang2024VideoHallucer hallucination detection in video 1800 Y/N N/A 12
CV-Benchtong2024CV-Bench vision-centric tasks 2638 open GPT
MM-SpuBenchye2024MM-SpuBench spurious biases in MLLMs 10773 MQA N/A 15
Charxivwang2024CharXiv diverse charts understanding 10k open GPT 24
MMRoli2024MMRo robot applications 26175 MQA, open GPT 13
Web2Codeyun2024Web2Code understanding of the web content 5990 Y/N, open GPT 4
MIA-Benchqian2024MIA-Bench layered instructions following 400 open GPT 29
CRABxu2024CRAB cross-environment tasks 100 N/A N/A 4
ScanReasonzhu2024ScanReason 3D reasoning grounding 10k bounding box N/A 11
MATHCHECK-GEOzhou2024MATHCHECK geometry reasoning 1440 open GPT 11
CHOPINLLMfan2024CHOPINLLM diverse charts understanding 6k open
MMIUmeng2024MMIU multi-image tasks 11K MQA N/A 24
GMAI-MMBenchchen2024GMAI-MMBench medical applications 26K MQA GPT 50
Besides, Zhu et al. zhu2024ScanReason introduced a benchmark named ScanReason, which required MLLMs to conduct joint reasoning on the question and the 3D environment before predicting the 3D locations of target objects. To evaluate MLLMs’ capability to accurately perceive fine-grained spatial relations such as depth, direction, and distance, Cheng et al. cheng2024SpatialRGPT proposed SpatialRGPT, which specifically focused on MLLMs’ ability to understand 3D spatial concepts like metric distance or size differences between objects. Meanwhile, a comprehensive evaluation benchmark for accurately assessing the capability of MLLMs on 3D tasks was crucial. In view of this, Li et al. li2023M3DBench introduced a comprehensive 3D-centric benchmark called M3DBench, which served as the foundation for developing a versatile and practical general-purpose assistant in real-world 3D environments.
7.4 Omnimodal
In the real world, people are surrounded by audio, images, videos, and text messages in their daily lives. These diverse modalities collectively enhance our ability to perceive and understand scenes. However, most benchmarks focus on a single modality, which limits their ability to adapt to real-world environments that involve multiple modalities. Therefore, some works focused on evaluating MLLMs’ capabilities in handling multiple modalities simultaneously. MusicAVQA li2022MusicAVQA was constructed to evaluate MLLMs’ ability to answer questions regarding visual objects, sounds, and their associations together. However, the visual scenes in MusicAVQA were limited to music performances, where the questions were only about instrument relationships, lacking exploration of more real-life scenarios. In response to this limitation, Yang et al. AVQA proposed AVQA, which was designed for audio-visual question answering on general videos of real-life scenarios. Furthermore, to cover a wider range of modal information, Chen et al. chen2024MCUB introduced a benchmark called the Multimodal Commonality Understanding Benchmark (MCUB), which included four modalities—image, audio, video, and point cloud. The task of MCUB was to measure the model’s ability to identify commonalities among input entities from diverse modalities, but overlooked the various other relationships among different modalities. To achieve general purpose multimodal intelligence, Ying et al.ying2024MMT-Bench present MMT-Bench, a comprehensive benchmark designed to assess MLLMs across massive multimodal tasks, which required expert knowledge and deliberate visual recognition, localization, reasoning, and planning.
8 Conclusion
Evaluation carries profound significance and is becoming imperative in the advancement of AGI models. It ensures that the models are not only performing as expected but also meeting the desired standards of accuracy, robustness, and fairness. Through rigorous evaluation, we can identify strengths and weaknesses, guide further improvements, and build trust in the deployment of AI systems in real-world applications. In this study, we provide a comprehensive overview of the evaluation and benchmarks of MLLMs, categorizing them into perception and understanding, cognition and reasoning, specific domains, key capabilities, and other modalities. We aim to enhance the understanding of the current status of MLLMs, elucidate their strengths and limitations, and provide insights into the future progression of MLLMs. Given the dynamic nature of this field, it’s possible that some recent developments may not be fully covered. To address this, we plan to continuously update and enhance the information on our website, incorporating new insights as they emerge.